OpenStudy (anonymous):

PLEASE HELP ME!!! A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $41,328. The variable costs will be $8.75 per book. The publisher will sell the finished product to bookstores at a price of $19 per book.

OpenStudy (anonymous):

I LOVE BOOBIES!!! :ddd

OpenStudy (anonymous):

12,056

OpenStudy (anonymous):

12,056

OpenStudy (anonymous):

Well i have 4032

OpenStudy (anonymous):

41328+8.75x=19x x= 4032 books

OpenStudy (anonymous):

i got 4,032, diogo is right

OpenStudy (anonymous):

15 + 0.09x = 8 + 0.14x

OpenStudy (anonymous):

x= time in minutes.

OpenStudy (anonymous):

140

OpenStudy (anonymous):

Yes, it is :)

OpenStudy (anonymous):

I guess so.. But why dont you try to understand the logic behind it instead of copying the final answer? Its so simple. Everytime you have a fixed cost and a variable cost, you need to put them in the same equation member. Since you want to find the interception between 2 given options, you just need to put "one = the other". Post some more and try to do it on your own. I'll help you

OpenStudy (anonymous):

Nice. Keep them comming then. I'll help you

OpenStudy (anonymous):

Is this the only information you have?

OpenStudy (anonymous):

So in this case you have a 2 system of equations. You need to separate the pounds and the price so it will be something like this: 1 pound * the amount of A + 1 pound * the amount of B = 134 pounds AND 4.35 dollars * the amount of A + 5.85 dollars * the amount of B = 699.9 dollars OR: 4.35A+5.85B=699.9 A+B=134 By solving this you'll get the amount of every product. A = 56 B = 78

OpenStudy (anonymous):

18

OpenStudy (anonymous):

I'll read the question real quick

OpenStudy (anonymous):

what can i asset you with?

OpenStudy (anonymous):

type A coffee would be 56 but idk about type B

OpenStudy (anonymous):

yush

OpenStudy (anonymous):

or at least think so

OpenStudy (anonymous):

yes ax is right it would be 56 for A and 78 for B

OpenStudy (anonymous):

oh this should be easy

OpenStudy (anonymous):

wait oh hmmm let me think again

OpenStudy (anonymous):

hmmm I think I got it

OpenStudy (anonymous):

hmm I think its 44

OpenStudy (anonymous):

@Fransisco_w please help mee

OpenStudy (anonymous):

dang well i estimated I actually got 43.5

OpenStudy (anonymous):

@Axellevi what do you think?

OpenStudy (anonymous):

i has no answer :(

OpenStudy (anonymous):

i sowwiee

OpenStudy (anonymous):

dang sorry

OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.4 megabytes (MB). The size of the high-quality version is 4.1 MB. Yesterday, the high quality version was downloaded twice as often as the standard version. The total size downloaded for the two versions was 2650 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

can you help me with this one @UltimateWitherLord

OpenStudy (anonymous):

@Axellevi help me with this one please

OpenStudy (anonymous):

hmmmmmmmmm im thinking

OpenStudy (anonymous):

number of standard version downloads?

OpenStudy (anonymous):

I cant think of the answer

OpenStudy (anonymous):

i cant help you with this one sorry